OpenStudy (anonymous):

which of the following is the least profitable investment for a game shop earning $2 profit from every game sold? A.Vehicle with $120 daily operating cost, delivering 80 games per day B.Inventory software with $10 daily operating cost, saving 30 games per day from theft C.Game cleaning system with $5 daily operating cost, increasing sales by 15 games per day D. electronic billboard with $20 daily operating cost, increasing sales by 20 games per day

OpenStudy (anonymous):

I think its D

OpenStudy (anonymous):

Here is my reasoning: I set up equations of the form Profit -Operating costs A) 80(2)-120=40 B) 30(2)-10=50 C) 15(2)-5 = 25 D)20(2)-20=20 The last option you make the least money soo that would be my preferred guess